A conglomerate corporation? develops, manufactures, and markets a wide range of? products, including medical diagnostic imaging? devices, jet? engines, lighting?products, and chemicals. In? 2013, the stock price rose

32.3632.36?%,

and in? 2014, the stock price declined

10.210.2?%.

If one purchased? $1,000 of some social media stock at the start of? 2013, its value would be

?$2 comma 953.042,953.04

at the end of 2014. Complete parts? (a) through? (c) below.

a. Compute the geometric mean rate of return per year of the? two-year period

2013dash–2014.

?(Hint: Denote an increase of

30.7630.76?%

as

Upper R 1R1equals=0.30760.3076?).

The geometric mean rate of return is per year.

b. If you purchased? $1,000 of this conglomerate? corporation's stock at the start of? 2013, what was its value at the end of? 2014?

The value of this stock at the end of 2014 was

?$

c. Compare the result of? (b) to the value of the? $1,000 of the social media stock. Choose the correct answer below and fill in the answer box to complete your choice.
Answer #1

a. Geometric mean rate of return is calculated using the following formula:

GR = \sqrt{1.3236*0.898} - 1

Geometric rate of return = 9%

b. If $1000 stock is purchased at the start of 2013:

Value in the end of 2013 = 1.3236*1000 = $1323

Value in the end of 2014 = 0.898*1323 = $1189

c. The stock value has increased by 18.9% from the year 2013-14. It has increased from $1000 to $1189.

In the question, there are no options. Please upload the question again if you don't find the correct option from this answer.